# Torque-based Graph Surgery: Enhancing Graph Neural Networks with Hierarchical Rewiring

### v1

@@ 4.2 Adjust Message Passing
The final node representations from the last layer are fed into a fully connected layer parameterized by $\mathbf{\Phi}\in\mathbb{R}^{m\times c}$ to yield the predicted class probabilities. These predictions are compared against the ground-truth labels using a cross-entropy loss minimized through gradient-based optimization.
4.3 Energy Calibration Objective
The energy function represents the unnormalized negative log-likelihood of node $v_{i}$, where a lower energy corresponds to a higher likelihood and, consequently, a greater probability of the sample being benign. In EBMs, the fundamental objective is to assign low energy to positive samples and high energy to others, thereby maximizing the likelihood of the observed data. This is typically achieved by minimizing the negative log-likelihood, formulated as
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\text{log}p_{\mathbf{\Theta}}(\mathbf{x}_{i}).$ (13)
[…]
$\begin{split}\mathcal{F}\big{(}p_{data}(\mathbf{x})\|p_{\mathbf{\Theta}}(\mathbf{x})\big{)}=\mathbb{E}_{p_{data}(\mathbf{x})}\Big{[}\frac{1}{2}\|\nabla_{\mathbf{\mathbf{x}}}\log p_{data}(\mathbf{x})-\nabla_{\mathbf{x}}\log p_{\mathbf{\Theta}}(\mathbf{x})\|^{2}\Big{]}.\end{split}$ (14)
To enlarge the energy signal of undesirable connections to facilitate more clearer distinction during edge selection, inspired by [46], we generate negative examples by randomly perturbing the graph structure, i.e., by adding or removing edges. Let $\widehat{E}_{\mathbf{\Theta}}(\mathbf{x}_{i})$ denote the energy of node $v_{i}$ on this perturbed graph. We then define an energy-based score according to the following definition
Definition 1
(Energy Score): For a node $v_{i}$, its score $S_{\mathbf{\Theta}}(\mathbf{x}_{i})$ assigned by the energy model is surrogated by a gradient in the discrete space:
$S_{\mathbf{\Theta}}(\mathbf{x}_{i})=\nabla_{\mathbf{x}_{i}}E_{\mathbf{\Theta}}(\mathbf{x}_{i})=\Big{[}\frac{E_{\mathbf{\Theta}}(\mathbf{x}_{1})-\widehat{E}_{\mathbf{\Theta}}(\mathbf{x}_{1})}{E_{\mathbf{\Theta}}(\mathbf{x}_{1})},\cdots,\frac{E_{\mathbf{\Theta}}(\mathbf{x}_{M})-\widehat{E}_{\mathbf{\Theta}}(\mathbf{x}_{M})}{E_{\mathbf{\Theta}}(\mathbf{x}_{M})}\Big{]},$ (15)
where $M$ is the number of perturbed samples used. In fact, since $-\log p_{\mathbf{\Theta}}(\mathbf{x})=E_{\mathbf{\Theta}}(\mathbf{x})$, Eq. (15) equals to the following formula,
$\begin{split}S_{\mathbf{\Theta}}(\mathbf{x}_{i})=\nabla_{\mathbf{x}_{i}}E_{\mathbf{\Theta}}(\mathbf{x}_{i})=\Big{[}\frac{\log p_{\mathbf{\Theta}}(\mathbf{x}_{1})-\log\widehat{p}_{\mathbf{\Theta}}(\mathbf{x}_{1})}{\log p_{\mathbf{\Theta}}(\mathbf{x}_{1})},\cdots,\frac{\log p_{\mathbf{\Theta}}(\mathbf{x}_{M})-\log\widehat{p}_{\mathbf{\Theta}}(\mathbf{x}_{M})}{\log p_{\mathbf{\Theta}}(\mathbf{x}_{M})}\Big{]}.\end{split}$ (16)
Due to the energy score surrogate, the objective (14) of score matching is reformulated as follows
$\mathcal{F}\big{(}p_{data}(\mathbf{x}_{i})\|p_{\mathbf{\Theta}}(\mathbf{x}_{i})\big{)}=\mathbb{E}_{p_{data}(\mathbf{x}_{i})}\Big{[}\frac{1}{2}\|S_{data}(\mathbf{x}_{i})-S_{\mathbf{\Theta}}(\mathbf{x}_{i})\|^{2}\Big{]}.$ (17)
However, the real data is unavailable during training, and thus following [49, 46] and Definition 1, the final optimization formula is given by
$\mathcal{L}_{sm}=\frac{1}{N}\sum_{i=1}^{N}\Big{[}\nabla_{\mathbf{x}_{i}}S_{\mathbf{\Theta}}(\mathbf{x}_{i})^{\top}\nabla_{\mathbf{x}_{i}}S_{\mathbf{\Theta}}(\mathbf{x}_{i})+\frac{1}{2}\|S_{\mathbf{\Theta}}(\mathbf{x}_{i})\|^{2}\Big{]}.$ (18)
By minimizing the surrogate loss (18), we not only approximate the original score matching objective of the EBM, but also explicitly amplify the energy signals associated with undesirable edges.
@@ 5 Experiments
Datasets. We evaluate our method on ten standard node classification benchmarks, including three homophilous graphs (Citeseer, Cora, and Pubmed) and seven heterophilous graphs (Texas, Wisconsin, Cornell, Actor, Tolokers, Questions, and Penn94). Among them, Tolokers, Questions, and Penn94 are large-scale datasets. Following previous work, we adopt the same data split strategy: 48% of the nodes are used for training, 32% for validation, and the remaining 20% for testing. The statistics of these datasets are summarized in Table 1, and further details can be found in Appendix C.2.
@@
Citeseer 0.74 3,327 4,676 7 3,703
Cora 0.81 2,708 5,278 6 1,433
Pubmed 0.80 19,717 44,327 3 500
Competitors. To evaluate the effectiveness of our model in capturing high-order interactions, we adopt APPNP as the downstream GNN, due to its suitability for deep graph architectures. We compare our approach against eleven baselines: a two-layer MLP; three classic homophily-oriented GNNs, including GCN [50], GAT [41] and APPNP [47]; and seven representative methods specifically designed for heterophilous graphs: H2GCN [51], GPR-GNN [52], FAGCN [39], ACM-GNN [38], OGNN [53], M2M-GNN [40], PCNet [54]. Furthermore, to assess the impact of our THR rewiring strategy, we compare it with other rewiring strategies: first-order spectral rewiring (FoSR) [55], joint denoising and rewiring (JDR) [36] and randomly removing edges (DropEdge) [56]. More details on all methods are provided in Appendix C.3.
Setups. We report node classification ACC, defined as the proportion of correctly predicted labels. For most benchmark datasets, models are trained using the Adam optimizer for 100 epochs with a learning rate of 0.001 and 512 hidden units. For larger datasets such as Tolokers, Questions, and Penn94, the number of training epochs is increased to 500 to ensure sufficient training. Detailed hyperparameter and environment configurations are provided in Appendix C.4, including the code link in C.1. Each experiment is repeated ten times, and we report the mean and standard deviation.
Node Classification Results. Table 2 reports test-set accuracy for three TorqueGNN variants, including edge-addition THR (A-THR), edge-removal THR (R-THR) and mixed THR (M-THR), alongside eleven state-of-the-art baselines across seven benchmark datasets. Several insights emerge: 1) On every dataset, at least one TorqueGNN variant appears among the top three performers, and TorqueGNN achieves the highest accuracy on the majority of benchmarks. 2) Three rewiring strategies yield substantial gains over APPNP and improvements are most pronounced on heterophilous graphs, underscoring the significance of our torque-based rewiring. 3) On Citeseer, Cora, and Pubmed, both A-THR and R-THR outperform the M-THR variant, suggesting that simultaneous addition and removal can inadvertently constrain information flow for citation networks. 4) In highly heterophilous datasets, although edge addition boosts APPNP, original undesirable links limit its effect. By contrast, R-THR achieves more high accuracy by judiciously deleting these connections. 5) In cases such as Pubmed and Cornell, R-THR attains the best results, likely because aggressive removal of noisy original edges without introducing additional candidates yields a cleaner propagation structure.
@@
Table 2: Node classification results on benchmark datasets: Mean ACC % ± Standard Deviation %. he first-, second- and third-best accuracy are highlighted in red, green and blue, respectively.
@@
Methods Citeseer Cora Pubmed Texas Wiscon. Cornell Actor
MLP 74.02±1.90 75.69±2.00 87.16±0.37 80.81±4.75 85.29±3.31 81.89±6.40 36.53±0.70
GCN 76.50±1.36 86.98±1.27 88.42±0.50 55.14±5.16 51.76±3.06 60.54±5.30 27.32±1.10
GAT 76.55±1.23 87.30±1.10 86.33±0.48 52.16±6.63 49.41±4.09 61.89±5.05 27.44±0.89
H2GCN 77.11±1.57 87.87±1.20 89.49±0.38 84.86±7.23 87.65±4.98 82.70±5.28 35.70±1.00
GPR-GNN 77.13±1.67 87.95±1.18 87.54±0.38 78.38±4.36 82.94±4.21 80.27±8.11 34.63±1.22
FAGCN 74.01±1.85 86.34±0.67 76.57±1.88 77.56±6.11 79.41±6.55 78.64±5.47 34.85±1.61
ACM-GCN 77.32±1.70 87.91±0.95 90.00±0.52 87.84±4.40 88.43±3.22 85.14±6.07 36.28±1.09
OGNN 77.31±1.73 88.37±0.75 90.15±0.38 86.22±4.12 88.04±3.63 87.03±4.73 37.99±1.00
M2M-GNN 77.20±1.80 88.12±1.00 90.35±0.60 89.19±4.50 89.01±4.10 86.48±6.10 36.72±1.60
PCNet 77.50±1.06 88.41±0.66 89.51±0.28 88.11±2.17 88.63±2.75 82.61±2.70 37.80±0.64
APPNP (backbone) 74.02±0.38 88.32±0.41 84.53±0.34 49.17±3.30 47.60±4.54 55.83±4.02 29.03±1.12
TorqueGNN w M-THR 75.86±0.23 84.78±0.28 87.52±0.06 91.11±1.67 89.60±1.20 89.44±1.67 36.72±0.16
TorqueGNN w R-THR 76.51±0.09 86.73±0.50 90.18±0.24 88.61±0.83 86.20±0.60 90.83±2.50 36.07±0.36
TorqueGNN w A-THR 78.03±0.05 88.47±0.28 89.38±0.02 86.11±0.01 82.00±0.01 83.61±3.15 36.24±0.04
@@
Table 3: Node classification results on large-scale datasets: Mean ACC % (ROC AUC for imbalanced Questions) ± Standard Deviation %, where the optimal and suboptimal results are highlighted in red and green, respectively.
Penn94 Tolokers Questions
baseGNN 75.19 (0.17) 76.47 (1.16) 46.57 (0.71)
FoSR 75.21 (0.19) 76.64 (1.15) 46.95 (0.83)
JDR 75.19 (0.23) 78.06 (0.88) 46.98 (0.87)
DropEdge 78.46 (0.02) 77.03 (0.26) 60.01 (0.24)
THR 80.88 (0.01) 78.08 (0.16) 62.64 (0.89)
Results on Larger Graphs. Scalability of rewiring techniques on large graphs is crucial, particularly for end-to-end methods that dynamically add and remove edges during training. In TorqueGNN, the principal computational burden stems from computing high-order torques and corresponding gaps, which incurs a complexity of $\mathcal{O}(|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$ (see Appendix B for details). Despite this overhead, THR remains tractable for large graphs. Table 3 compares multiple rewiring schemes on larger heterophilous datasets: Both THR and DropEdge yield substantial accuracy improvements over the base GNN, yet THR consistently outperforms all alternatives. In contrast, FoSR and JDR achieve only marginal gains. Moreover, THR surpasses DropEdge, validating the effectiveness of the proposed torque-driven manner.
Figure 2: Results of TorqueGNN w/o THR and w THR on graphs with noisy edges.
Results on Adversarially Perturbed Graphs. Our torque-driven rewiring not only improves predictive accuracy but also enhances robustness to structural attacks by detecting and suppressing adversarial edges. To evaluate this, we employ MetaAttack [57] to inject adversarial links at a perturbation rate of 25% of the original edge count. Figure 2 compares classification accuracy under three settings: TorqueGNN without THR (w/o THR) on attacked graph, TorqueGNN with THR (w THR) on attacked graph and TorqueGNN w THR on normal graph. In all cases, we employ the mixed addition and deletion strategy for optimization. As shown, across both homophilous and heterophilous datasets, applying THR to the perturbed graph substantially improves classification performance, which demonstrates the effectiveness of our method in detecting and eliminating adversarial edges. Moreover, on heterophilous datasets, TorqueGNN w THR applied to the attacked graph even surpasses its performance on the normal graph, suggesting that the injected adversarial edges accentuate the distinction between desirable and undesirable connections, thereby enhancing the effectiveness of our rewiring mechanism.
@@
Figure 3: Ablation study: Performance comparison of GCN and GAT with different rewiring schemes on four datasets.
Ablation Study. Table 2 has shown that all three THR variants markedly improve upon the APPNP backbone. To evaluate their broader applicability, we incorporate each strategy into two classic two-layer GNNs (GCN and GAT), and repeat the ablation studies in Figure 3. As illustrated, on the highly heterophilous Texas, Wisconsin, and Actor datasets, every rewiring variant substantially outperforms its base GNN. Even on the homophilous Pubmed, R-THR surpasses the base models, echoing the results in Table 2 and underscoring the broad applicability of our torque-driven optimization. We note, however, that A-THR and M-THR occasionally underperform relative to the baselines, likely because the current structure is already well-formed, and excessive edge additions may introduce noise, thereby degrading classification performance. Ablation study of significant components is included in Appendix C.5.
Parameter Analysis. To assess the influence of the key hyperparameter $\alpha$, we perform a sensitivity analysis shown in Figure 4. The figure examine the trade-off parameter $\alpha$, which balances the contribution of the learned hidden representation and the original input features, where a larger $\alpha$ increases the influence of the hidden representations. We observe that smaller graphs (e.g., Texas and Cornell), optimal accuracy is achieved at low $\alpha$ value ($0.01\leq\alpha\leq 0.05$), implying that raw node features provide sufficient discriminative power. In contrast, on medium-to-large graphs, better performance is observed when $\alpha\geq 0.5$, reflecting the necessity of high-order hidden representations to capture more complex community structures. Sensitivity analysis of other hyperparameters is presented in Appendix C.5.

### v3

@@ 4.2 Adjust Message Passing
Here, $\alpha$ controls the trade-off between the hidden representation and the residual connection. The initial representation $\mathbf{h}_{i}^{(0)}=\mathbf{x}_{i}\mathbf{\Theta}$ is computed through a linear transformation of the input feature $\mathbf{x}_{i}$. The final node representations from the last layer are passed through a fully connected layer parameterized by $\mathbf{\Phi}\in\mathbb{R}^{m\times c}$, which yields the predicted class probabilities. These predictions are compared against the ground-truth labels using a cross-entropy loss, which is minimized through gradient-based optimization.
5 Complexity Analysis
The dominant computational cost of THR lies in: 1) Torque computation and graph rewiring. For each order $l$, we compute torque values only on the edges in $\mathcal{A}^{(l)}$, costing $\mathcal{O}(|\mathcal{A}^{(l)}|)$, and then sort these values in $\mathcal{O}(|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$. When adding edges, if the candidate set size is $B$, the combined probability calculation and sorting cost is $\mathcal{O}(B+B\log B)$. 2) Message passing on the rewired graph $\mathcal{A}^{(l)}$. For the input layer with parameter $\mathbf{\Theta}\in\mathbb{R}^{d\times m}$ on $\mathbf{X}\in\mathbb{R}^{N\times d}$, it costs $\mathcal{O}(Ndm)$. Aggregation over $\mathcal{A}^{(l)}$ then costs $\mathcal{O}(m|\mathcal{A}^{(l)}|)$ per layer. The output layer with $\mathbf{\Phi}\in\mathbb{R}^{m\times c}$ requires $\mathcal{O}(Nmc)$. Putting these together for an $L$-layer network and assuming $B\ll|\mathcal{A}^{(l)}|$ for all $l$, the overall complexity is $\mathcal{O}(Ndm+\sum_{l=1}^{L}|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$, which is slightly higher than that of standard methods with $\mathcal{O}(Ndm+m|\mathcal{A}^{(l)}|)$.
@@ 6 Experiments
Datasets. We evaluate our method on eleven standard node classification benchmarks, which include six heterophilic datasets: Texas, Wisconsin, Cornell, Actor, Penn94 and Flickr; five homophilous graphs Citeseer, Cora, Pubmed, Tolokers and Questions. Among them, Tolokers, Questions, Penn94 and Flickr are large-scale datasets. The statistics for these datasets are summarized in Table 1, with further details provided in Appendix C.2.
@@
Flickr 0.32 89,250 2,724,458 7 500
Baselines. THR is a plug-in module that can be integrated into various state-of-the-art GNNs. To evaluate the improvements offered by THR for GNNs, we select three representative models for experimentation, including two models designed for homophilous graphs: the vanilla GCN [45] and the deep-based APPNP [43], and GPRGNN [46] that is designed for heterophilous graphs.
To evaluate the effectiveness of THR in comparison to other graph rewiring techniques, we select five superior methods, including: First-order Dpectral Rewiring (FoSR) [32], Batch Ollivier-Ricci Flow (BORF) [47], Stochastic Jost and Liu Curvature Rewiring (SJLR) [48], Deep Heterophily Graph Rewiring (DHGR) [14] and randomly edge removal (DropEdge). Here, we adopt layer-wise DropEdge (Dropedge-L), as proposed by [49], to ensure a fair comparison with the hierarchical structure of THR. Further details on all methods are provided in Appendix C.3.
Setups. We report node classification accuracy (ACC), defined as the proportion of correctly predicted labels. For all benchmark datasets, models are trained using the Adam optimizer. Competitors are performed based on their respective source code. Detailed hyperparameters and environment configurations for THR are provided in Appendix C.4. Following prior work [37, 50], we adopt the data split strategy for all methods: 48% of the nodes are used for training, 32% for validation, and the remaining 20% for testing. Each experiment is conducted over 10 runs with different random splits, and the results are reported as the mean and standard deviation.
@@
Table 2: Node classification results on benchmark datasets with GCN and GPRGNN as the backbone models: Mean ACC % (Standard Deviation %). The first- and second-best accuracies are highlighted in red and green, respectively.
@@
Methods/Datasets Citeseer Cora Pubmed Texas Wisconsin Actor Cornell
GCN 75.52 (2.19) 86.96 (1.27) 86.43 (0.38) 58.61 (7.18) 52.60 (8.72) 30.15 (1.03) 57.50 (4.66)
FoSR 78.03 (1.45) 87.00 (1.21) 86.34 (0.31) 74.70 (6.23) 65.58 (4.89) 30.16 (1.03) 54.59 (5.01)
BROF 78.45 (1.52) 86.86 (1.35) 86.42 (0.38) 74.51 (6.26) 65.59 (4.52) 30.20 (1.17) 60.27 (3.64)
SJLR 77.87 (1.81) 86.60 (1.64) 86.52 (1.73) 60.14 (0.89) 55.16 (0.95) 30.80 (1.34) 58.11 (6.86)
DHGR 78.68 (1.51) 86.61 (1.73) 86.40 (0.38) 60.20 (6.39) 66.07 (12.51) 34.39 (0.99) 58.68 (5.01)
DropEdge-L 74.93 (1.85) 86.62 (1.23) 83.07 (2.58) 62.74 (8.32) 58.82 (8.24) 32.97 (0.92) 54.32 (3.72)
THR 80.43 (1.52) 86.97 (1.19) 87.21 (0.45) 76.27 (4.67) 68.09 (2.71) 33.20 (0.90) 58.91 (9.11)
GPRGNN 77.37 (1.83) 87.34 (1.14) 87.21 (0.43) 89.22 (5.56) 87.94 (5.29) 37.27 (1.16) 80.27 (6.63)
FoSR 77.37 (1.83) 87.52 (1.63) 87.22 (0.46) 90.20 (5.04) 89.85 (3.45) 37.25 (1.19) 84.05 (7.88)
BORF 78.77 (1.67) 87.49 (1.24) 87.17 (0.39) 91.16 (5.15) 89.11 (4.32) 37.52 (1.06) 85.49 (4.83)
SJLR 78.38 (1.49) 86.97 (1.63) 88.11 (0.41) 90.00 (2.83) 89.26 (6.38) 34.87 (1.69) 81.62 (9.35)
DHGR 77.77 (2.06) 87.19 (1.39) 87.69 (0.47) 89.02 (4.31) 86.03 (6.32) 35.20 (1.20) 84.31 (4.56)
DropEdge-L 78.73 (1.91) 86.91 (1.07) 87.50 (0.48) 90.17 (3.06) 87.79 (6.28) 37.77 (1.16) 84.05 (9.00)
THR 79.15 (1.69) 87.60 (1.15) 88.28 (0.52) 91.96 (3.76) 91.91 (4.75) 38.00 (0.56) 86.22 (5.19)
Node Classification Results. Table 2 presents the test-set accuracy gains achieved by various rewiring approaches on GCN and GPRGNN across seven benchmark datasets. Several key insights can be drawn: 1) Compared to the baselines, all rewiring methods show performance improvements on most datasets, with particularly notable gains on heterophilous graphs. 2) In all datasets, the proposed THR ranks among the top two performers, achieving the highest accuracy gain on the majority of benchmarks. 3) Although FoSR, BORF, and DHGR also exhibit strong performance on certain datasets, their gains are only marginally higher than those of THR. Overall, THR outperforms these methods and delivers the best results in all cases when GPRGNN is used as the downstream model. 4) DropEdge-L, which is also based on hierarchical graph rewiring, outperforms other rewiring methods on some datasets (e.g., Texas and Actor), validating the effectiveness of the hierarchical strategy. Although DropEdge shows performance improvements on certain datasets, its inherent randomness negatively impacts the model’s performance, resulting in lower performance than the baseline in some cases, e.g., Citeseer. This further validates the effectiveness of the proposed torque-driven hierarchical approach.
@@
Table 3: Node classification results on large-scale datasets: Mean ACC % (ROC AUC for imbalanced Questions and Tolokers) (Standard Deviation %), where the optimal and suboptimal results are highlighted in red and green, respectively. OoM means that the model suffers from the out-of-memory error.
Methods/Datasets Questions Tolokers Penn94 Flickr
GCN 75.26 (0.84) 83.79 (0.74) 80.18 (0.36) 57.48 (7.85)
FoSR 75.19 (0.71) 84.14 (0.99) 80.19 (0.35) 58.03 (6.75)
BROF 75.15 (0.84) MemoryError OoM OoM
SJLR 72.07 (6.12) 76.14 (1.14) 80.20 (0.28) 64.49 (2.82)
DHGR OoM 76.45 (12.16) OoM OoM
DropEdge-L 74.06 (1.11) 84.00 (0.65) 62.27 (0.35) 59.29 (2.26)
THR 75.92 (1.09) 84.43 (0.88) 80.32 (0.23) 68.29 (0.78)
GPRGNN 72.89 (1.42) 71.99 (0.93) 84.18 (0.30) 48.90 (5.81)
FoSR 72.91 (1.43) 71.99 (0.93) 84.22 (0.29) 49. 16 (6.49)
BORF 72.99 (1.44) MemoryError OoM OoM
SJLR 72.27 (1.24) 69.46 (1.07) 83.89 (0.20) 61.61 (3.22)
DHGR OoM 70.96 (1.14) OoM OoM
DropEdge-L 72.07 (1.35) 71.98 (1.09) 83.67 (0.44) 63.42 (3.26)
THR 73.41 (0.98) 72.05 (1.24) 84.45 (0.29) 65.29 (3.30)
Results on Larger Graphs. Scalability of rewiring techniques on large graphs is crucial, particularly for end-to-end methods that dynamically add and remove edges during training. In THR, the primary computational cost arises from computing torques and the corresponding gaps, which incurs a complexity of $\mathcal{O}(|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$ (see Section 5 for details). Despite this overhead, THR remains computationally feasible for large graphs. Table 3 compares several rewiring schemes on larger datasets, with THR consistently outperforming all alternatives. Notably, with the exception of the Flickr dataset, all rewiring methods show only marginal improvements, and in some cases, even lead to a decline in performance. This may be attributed to the fact that the raw graphs of these datasets already contain sufficient structural information, and the rewiring methods introduce only minor modifications, or in some cases, may even result in the loss of critical semantics, thus negatively impacting classification performance.
@@
Figure 2: Ablation study: Performance comparison of GCN, GPRGNN, and APPNP with various THR variants across six datasets.
Ablation Study. We conduct an ablation study to assess the impact of edge removal and addition operations in THR, using GCN, GPRGNN, and APPNP as backbone models. THR has three variants: edge-addition THR (A-THR), edge-removal THR (R-THR), and mixed THR (M-THR). As illustrated in Figure 2, most rewiring variants significantly outperform their base GNNs. However, on the Wisconsin dataset, GPRGNN slightly surpasses GPRGNN w R-THR, likely because GPRGNN effectively allocates sign edges to distinguish class information, while R-THR removes heterophilic links, inadvertently causing the model to lose some discriminative features. On the Flickr dataset, R-THR improves performance for all models, while A-THR and M-THR degrade the performance of APPNP. Similarly, on PubMed, both A-THR and M-THR reduce the performance of GCN. These results suggest that for these graphs, excessive edge addition leads to information interference and confusion of node features, while GPRGNN mitigates this effect by utilizing a sign edge strategy. In conclusion, the THR strategy enhances model performance, but its effectiveness varies across datasets with different characteristics.
@@
Table 4: Ablation study: A comparison of THR and its variants by removing specific components. The optimal and suboptimal results are highlighted in bold and underlined, respectively.
Datasets Texas Wisconsin Cornell Actor Penn94 Flickr
[…]
THR 72.01 (6.13) 75.89 (3.46) 77.30 (7.17) 36.28 (1.15) 76.21 (0.47) 63.28 (1.56)
Parameter Analysis. Since the edge-removal procedure automatically determines the cutoff $\mathcal{K}$, we investigate the main hyperparameter $t$ of THR, which defines the number of candidate edges for addition. As shown in Figure 3, we present the performance curves for varying $t$ values in $\{2,4,6,8,10\}$ across five datasets. On both homophilous and heterophilous datasets, accuracy increases as $t$ grows, demonstrating that the proposed edge-addition scheme aids the model in capturing global information. However, this does not imply that adding more edges is always beneficial. For instance, on the Flickr dataset, performance decreases when $t=8$, as excessive edge addition may introduce noise, as highlighted in the ablation study. Sensitivity analysis of other hyperparameters is presented in Appendix C.5.
Figure 3: Parameter sensitivity: Performance curves on five datasets as the number of candidate edges $t$ varies from 2 to 10.